How much energy is required to raise the temperature of 18.6 g of water from 15°C to 134°C?
1
Expert's answer
2020-04-06T11:46:01-0400

The minimum energy required to raise the temperature of water can be calculated using the specific heat of water "c_w" , equal to 4.186 J/g/°C:

"Q = c_wm_w(T_2 -T_1) = 4.186\cdot18.6\cdot(134-15) = 9265" J, or 9.27 kJ

Answer: 9.27 kJ are needed to raise the temperature of 18.6 g of water from 15°C to 134°C.